### v1

# Products of hyperbolic spaces

Abstract
The class of uniformly smooth hyperbolic spaces was recently introduced by Pinto as a common generalization of both CAT(0) spaces and uniformly smooth Banach spaces, in a way that Reich’s theorem on resolvent convergence could still be proven. We define products of such spaces, showing that they are reasonably well-behaved. In this manner, we provide the first example of a space for which Reich’s theorem holds and which is neither a CAT(0) space, nor a convex subset of a normed space.
Mathematics Subject Classification 2020: 51F99, 52A55, 53C23.
Keywords: Product spaces, uniform convexity, uniform smoothness, hyperbolic spaces.
1 Introduction
In the last few decades, there has been a renewed interest in nonlinear (metric) generalizations of the classical structures of functional analysis, primarily motivated by fixed point theory and convex optimization. The center of these developments has been the class of CAT(0) spaces, which are metric spaces that aim to model the idea of non-positive curvature (directly generalizing Hilbert spaces). More general spaces of a ‘hyperbolic’ nature have also been studied, and a particularly flexible notion has been introduced by Kohlenbach [9] under the name of $W$-hyperbolic spaces (see the next section for a definition and [10, pp. 384–387] for a detailed discussion on the relationship between various definitions of hyperbolicity).
Recently, Pinto [18] has defined a subclass of $W$-hyperbolic spaces which he called uniformly smooth hyperbolic spaces. He was motivated by more than just generalizing the Banach space notion of uniform smoothness, so we shall step back a bit in order to understand the context.
If $X$ is a metric space possessing enough convexity structure, and $T:X\to X$ is a nonexpansive mapping, one can define, for any $x\in X$, the point $x_{t}$, called a ‘resolvent’, by the implicit definition $x_{t}=tx_{t}+(1-t)x$ (existence and uniqueness being guaranteed by the Banach contraction principle). One may then ask about the strong convergence of the path $(x_{t})$ (for $t\to 1$), and, for $X$ being a bounded convex subset of a Hilbert space, this was first proven in 1967, independently by Browder [5] and Halpern [8]. A little more than ten years later, the corresponding result for uniformly smooth Banach spaces, requiring significantly new ideas, was proven by Reich [20].
In the past two decades, these kinds of convergence results have begun to be studied from the point of view of proof mining, which is a research paradigm (first suggested by Kreisel in the 1950s under the name of ‘unwinding of proofs’ and then highly developed starting in the 1990s by Kohlenbach and his students and collaborators) that aims to analyze ordinary mathematical proofs using tools from proof theory (a branch of mathematical logic) in order to extract additional – typically quantitative – information out of them (for more information, see Kohlenbach’s monograph [10] and his recent research survey [12]).
The analysis of the Browder-Halpern result was obtained by Kohlenbach in 2011 [11]; its sequel, the work on Reich’s theorem, had to wait, like the original theorem itself, a decade until it was finally achieved by Kohlenbach and the author [14], being one of the most intricate results ever achieved by the proof mining program (see also the author’s recent extension [23] to Reich’s original setting of accretive operators). The goals of these endeavours were to obtain quantitative results, namely so-called ‘rates of metastability’; what is more relevant here is that a proof-theoretic analysis usually strips down a proof to its bare essence, leaving it more suitable to generalization.
Such a generalization was needed, as per Pinto [18], since the theorem had also been proven, in the meantime, in the setting of CAT(0) spaces by Saejung [21] (and then proof-theoretically analyzed by Kohlenbach and Leuştean [13]), and, thus, one had to have a framework to encompass both CAT(0) spaces and uniformly smooth Banach spaces, this being achieved in the class of uniformly smooth hyperbolic spaces. In addition, what Kohlenbach and the author had done to tame the proof of Reich’s theorem and make it amenable to analysis and to generalization was to introduce the additional uniform convexity hypothesis on the Banach space (thus, still maintaining $L^{p}$ spaces as a particular case of the theorem), a notion which had already been formulated in the hyperbolic setting by Leuştean [16, 17] in the form of $UCW$-hyperbolic spaces. Thus, Pinto was able to show that Reich’s theorem and its consequences do hold for uniformly smooth $UCW$-hyperbolic spaces. The generalization of the proof-theoretic analysis itself was ultimately relegated to the forthcoming paper [19].
Now, this work raises the question of whether the introduction of this new abstract notion only provides a unified proof of the theorem for the two known cases or whether one has properly extended the class of spaces for which the theorem holds. This is the question that this paper aims to answer.
Towards that end, we take the existing construction of a product of metric spaces and show that it extends all the way to uniformly smooth hyperbolic spaces. Uniform convexity may possibly not be preserved under products, but we show that a specific property, which is enough for the proof of Reich’s theorem, is. Finally, we give an example of a product space for which we show that it cannot be either a CAT(0) space or a bounded subset of a normed space.
Section 2 recalls the main facts about the classes of spaces under consideration. Section 3 presents all the results on the product construction, including the concrete example mentioned above.
2 Nonlinear spaces
One says that a metric space $(X,d)$ is geodesic if for any two points $x$, $y\in X$ there is a geodesic that joins them, i.e. a mapping $\gamma:[0,1]\to X$ such that $\gamma(0)=x$, $\gamma(1)=y$ and for any $t$, $t^{\prime}\in[0,1]$ we have that
$d(\gamma(t),\gamma(t^{\prime}))=|t-t^{\prime}|d(x,y).$
Generally, geodesics need not be unique, and, thus, one is led to the following definition (which, as stated in the Introduction, is due to Kohlenbach [9]), which considers geodesics on a metric space as an additional structure (and not as a property), which is required to satisfy additional properties.
Definition 2.1.
A $W$-hyperbolic space is a triple $(X,d,W)$ where $(X,d)$ is a metric space and $W:X^{2}\times[0,1]\to X$ such that, for all $x$, $y$, $z$, $w\in X$ and $\lambda$, $\mu\in[0,1]$, we have that
  1. (W1)
    $d(z,W(x,y,\lambda))\leq(1-\lambda)d(z,x)+\lambda d(z,y)$;
  2. (W2)
    $d(W(x,y,\lambda),W(x,y,\mu)=|\lambda-\mu|d(x,y)$;
  3. (W3)
    $W(x,y,\lambda)=W(y,x,1-\lambda)$;
  4. (W4)
    $d(W(x,z,\lambda),W(y,w,\lambda))\leq(1-\lambda)d(x,y)+\lambda d(z,w)$.
Clearly, any normed space may be made into a $W$-hyperbolic space in a canonical way.
[…]
Proposition 2.2.
Let $(X,d,W)$ be a $W$-hyperbolic space and $x$, $y$, $a\in X$. Then
$d^{2}\left(\frac{x+y}{2},a\right)\leq\frac{d^{2}(x,a)+d^{2}(y,a)}{2}.$
Proof.
Using $(W1)$, we get that
$d^{2}\left(\frac{x+y}{2},a\right)\leq\left(\frac{d(x,a)+d(y,a)}{2}\right)^{2}\leq\frac{d^{2}(x,a)+d^{2}(y,a)}{2}.$
As per [9, 16], a particular nonlinear class of $W$-hyperbolic spaces is the one of CAT(0) spaces, introduced by A. Aleksandrov [1] and named as such by M. Gromov [7], which may be defined as those geodesic spaces $(X,d)$ such that, for any geodesic $\gamma:[0,1]\to X$ and for any $z\in X$ and $t\in[0,1]$ we have that
$d^{2}(z,\gamma(t))\leq(1-t)d^{2}(z,\gamma(0))+td^{2}(z,\gamma(1))-t(1-t)d^{2}(\gamma(0),\gamma(1)).$ (1)
Another well-known fact about CAT(0) spaces is that each such space $(X,d)$ is uniquely geodesic – that is, for any $x$, $y\in X$ there is a unique geodesic $\gamma:[0,1]\to X$ that joins them – this giving their unique $W$-hyperbolic space structure.
In particular, any inner product space is a CAT(0) space, and experience shows that (complete) CAT(0) spaces may be regarded as the rightful nonlinear generalization of Hilbert spaces.
In 2008, Berg and Nikolaev proved (see [3, Proposition 14]) that in any metric space $(X,d)$, the function $\langle\cdot,\cdot\rangle:X^{2}\times X^{2}\to\mathbb{R}$, defined, for any $x$, $y$, $u$, $v\in X$, denoting an ordered pair of points $(a,b)\in X^{2}$ by $\overrightarrow{ab}$, by
$\langle\overrightarrow{xy},\overrightarrow{uv}\rangle:=\frac{1}{2}(d^{2}(x,v)+d^{2}(y,u)-d^{2}(x,u)-d^{2}(y,v)),$
called the quasi-linearization function, is the unique one such that, for any $x$, $y$, $u$, $v$, $w\in X$, we have that:
  1. (i)
    $\langle\overrightarrow{xy},\overrightarrow{xy}\rangle=d^{2}(x,y)$;
  2. (ii)
    $\langle\overrightarrow{xy},\overrightarrow{uv}\rangle=\langle\overrightarrow{uv},\overrightarrow{xy}\rangle$;
  3. (iii)
    $\langle\overrightarrow{yx},\overrightarrow{uv}\rangle=-\langle\overrightarrow{xy},\overrightarrow{uv}\rangle$;
  4. (iv)
    $\langle\overrightarrow{xy},\overrightarrow{uv}\rangle+\langle\overrightarrow{xy},\overrightarrow{vw}\rangle=\langle\overrightarrow{xy},\overrightarrow{uw}\rangle$.
The inner product notation is justified by the fact that if $X$ is a (real) Hilbert space, for any $x$, $y$, $u$, $v\in X$,
$\langle\overrightarrow{xy},\overrightarrow{uv}\rangle=\langle x-y,u-v\rangle=\langle y-x,v-u\rangle$
The main result of [3], namely Theorem 1 of said paper, characterized CAT(0) spaces as being exactly those geodesic spaces $(X,d)$ such that the corresponding Cauchy-Schwarz inequality is satisfied, i.e. for any $x$, $y$, $u$, $v\in X$,
$\langle\overrightarrow{xy},\overrightarrow{uv}\rangle\leq d(x,y)d(u,v).$
A representative example of a (complete) CAT(0) space is the (hyperbolic) Poincaré upper half-plane model (see also [4]), having as the underlying set
:=$\mathbb{H}:=\{(x_{1},x_{2})\in\mathbb{R}^{2}\mid x_{2}>0\},$
where, given the function $\operatorname{arcosh}:[1,\infty)\to[0,\infty)$, where for every $t\in[1,\infty)$, $\operatorname{arcosh}t=\ln(t+\sqrt{t^{2}-1})$, the distance function is defined as follows: for any $x=(x_{1},x_{2})$, $y=(y_{1},y_{2})\in\mathbb{H}$, one sets
$d(x,y):=\operatorname{arcosh}\left(1+\frac{(y_{1}-x_{1})^{2}+(y_{2}-x_{2})^{2}}{2x_{2}y_{2}}\right).$ (2)
One proves, usually by considering this distance function as arising from a Riemannian metric, that geodesic lines of this space are of two types: for every $a\in\mathbb{R}$ and $r>0$, one has the semicircle
$\mathcal{C}_{a,r}=\{(x_{1},x_{2})\in\mathbb{H}\mid(x_{1}-a)^{2}+x_{2}^{2}=r^{2}\},$
while for every $a\in\mathbb{R}$, one has the ray
$\mathcal{R}_{a}=\{(x_{1},x_{2})\in\mathbb{H}\mid x_{1}=a\}.$
It may be then easily shown that for every two points there is exactly one geodesic segment that joins them. The general formula for the convex combination of two points is somewhat involved and we shall omit it, instead giving only the specialized formula for the midpoint, which shall be used later. For any $x=(x_{1},x_{2})$, $y=(y_{1},y_{2})\in\mathbb{H}$, we have that
$W\left(x,y,\frac{1}{2}\right)=\left(\frac{x_{1}y_{2}+x_{2}y_{1}}{x_{2}+y_{2}},\frac{\sqrt{x_{2}y_{2}}\cdot\sqrt{(x_{2}+y_{2})^{2}+(x_{1}-y_{1})^{2}}}{x_{2}+y_{2}}\right).$ (3)
We will now explore the class of uniformly smooth Banach spaces, in order to better understand the way it was generalized in [18] to uniformly smooth hyperbolic spaces.
Definition 2.3.
Let $X$ be a Banach space. We define the normalized duality mapping of $X$ to be the map $J:X\to 2^{X^{*}}$, defined, for all $x\in X$, by
$J(x):=\{x^{*}\in X^{*}\mid x^{*}(x)=\|x\|^{2},\ \|x^{*}\|=\|x\|\}.$
A Banach space $X$ is called smooth if for any $x\in X$ with $\|x\|=1$, we have that for any $y\in X$ with $\|y\|=1$, the limit
$\lim_{h\to 0}\frac{\|x+hy\|-\|x\|}{h}$ (4)
exists, and uniformly smooth if the limit is attained uniformly in $x$ and $y$. The condition of (not necessarily uniform) smoothness has been proven to be equivalent to the fact that the normalized duality mapping of the space, $J:X\to 2^{X^{*}}$, is single-valued – and we shall denote its unique section by $j:X\to X^{*}$. Therefore, for all $x\in E$, $j(x)(x)=\|x\|^{2}$ and $\|j(x)\|=\|x\|$. Hilbert spaces are smooth, and clearly $j(x)(y)$ is then simply $\langle y,x\rangle$, for any $x,y$ in the space. Because of this, we may consider the $j$ to be a generalized variant of the inner product, sharing some of its nice properties. One generally denotes, for all spaces $X$, all $x^{*}\in X^{*}$ and $y\in X$, $x^{*}(y)$ by $\langle y,x^{*}\rangle$. In addition, the homogeneity of $j$ – i.e. that for all $x\in X$ and $t\in\mathbb{R}$, $j(tx)=tj(x)$ – follows immediately from the definition of the duality mapping.
This view of $j$ as a generalization of the inner product, together with the way the quasi-linearization function was seen to be defined above, led Pinto [18] to the following definition.
Definition 2.4 ([18, Section 3]).
A smooth hyperbolic space is a quadruple $(X,d,W,\pi)$, where $(X,d,W)$ is a $W$-hyperbolic space and $\pi:X^{2}\times X^{2}\to\mathbb{R}$, such that, for any $x$, $y$, $u$, $v\in X$ (where, again, an ordered pair of points $(a,b)\in X^{2}$ is denoted by $\overrightarrow{ab}$):
  1. (P1)
    $\pi(\overrightarrow{xy},\overrightarrow{xy})=d^{2}(x,y)$;
  2. (P2)
    $\pi(\overrightarrow{xy},\overrightarrow{uv})=-\pi(\overrightarrow{yx},\overrightarrow{uv})=-\pi(\overrightarrow{xy},\overrightarrow{vu})$;
  3. (P3)
    $\pi(\overrightarrow{xy},\overrightarrow{uv})+\pi(\overrightarrow{yz},\overrightarrow{uv})=\pi(\overrightarrow{xz},\overrightarrow{uv})$;
  4. (P4)
    $\pi(\overrightarrow{xy},\overrightarrow{uv})\leq d(x,y)d(u,v)$;
  5. (P5)
    $d^{2}(W(x,y,\lambda),z)\leq(1-\lambda)^{2}d^{2}(x,z)+2\lambda\pi(\overrightarrow{yz},\overrightarrow{W(x,y,\lambda)z})$.
It is a classical result that, in uniformly smooth Banach spaces, the mapping $j$ is norm-to-norm uniformly continuous on bounded subsets. In the PhD thesis of Bénilan [2, p. 0.5, Proposition 0.3], it is shown that the norm-to-norm uniform continuity on bounded subsets of an arbitrary duality selection mapping is in fact equivalent to uniform smoothness. A more recent proof which uses ideas due to Giles [6] may be found in [15, Appendix A]. Motivated by this fact, Pinto defined a uniformly smooth hyperbolic space to be a smooth hyperbolic space $(X,d,W,\pi)$ such that there is an $\omega:(0,\infty)\times(0,\infty)\to(0,\infty)$, called a modulus of uniform continuity for $\pi$, having the property that, for any $r$, $\varepsilon>0$ and $a$, $u$, $v$, $x$, $y\in X$ with $d(u,a)\leq r$, $d(v,a)\leq r$ and $d(u,v)\leq\omega(r,\varepsilon)$, one has that
$|\pi(\overrightarrow{xy},\overrightarrow{ua})-\pi(\overrightarrow{xy},\overrightarrow{va})|\leq\varepsilon\cdot d(x,y).$
As mentioned in the Introduction, the remaining class of Banach spaces whose nonlinear generalization is relevant to our purposes is that of uniformly convex Banach spaces. We shall now turn to exploring uniform convexity directly in the hyperbolic setting, as first introduced by Leuştean [16, 17].
Definition 2.5.
If $(X,d,W)$ is a $W$-hyperbolic space, then a modulus of uniform convexity for $(X,d,W)$ is a function $\eta:(0,\infty)\times(0,\infty)\to(0,1]$ such that, for any $r$, $\varepsilon>0$ and any $a$, $x$, $y\in X$ with $d(x,a)\leq r$, $d(y,a)\leq r$, $d(x,y)\geq\varepsilon r$, we have that
$d\left(\frac{x+y}{2},a\right)\leq(1-\eta(r,\varepsilon))r.$
We call the modulus monotone if, for any $r$, $s$, $\varepsilon>0$ with $s\leq r$, we have $\eta(r,\varepsilon)\leq\eta(s,\varepsilon)$.
Definition 2.6.
A $UCW$-hyperbolic space is a $W$-hyperbolic space that admits a monotone modulus of uniform convexity.
As remarked in [16, Proposition 2.6], CAT(0) spaces are $UCW$-hyperbolic spaces having as a modulus of uniform convexity the function $(r,\varepsilon)\mapsto\frac{\varepsilon^{2}}{8}$, quadratic in $\varepsilon$.
The crucial property of uniform convexity that was used in [14] was the following, which we now codify for the first time.
Definition 2.7.
Let $\psi:(0,\infty)\times(0,\infty)\to(0,\infty)$. We say that a $W$-hyperbolic space $(X,d,W)$ has property $(G)$ with modulus $\psi$ if, for any $r$, $\varepsilon>0$ and any $a$, $x$, $y\in X$ with $d(x,a)\leq r$, $d(y,a)\leq r$, $d(x,y)\geq\varepsilon$, we have that
$d^{2}\left(\frac{x+y}{2},a\right)\leq\frac{1}{2}d^{2}(x,a)+\frac{1}{2}d^{2}(y,a)-\psi(r,\varepsilon).$
By [14, Proposition 2.4] (among others, see the references given there; this result essentially goes back to Zălinescu [24, Section 4]), uniformly convex Banach spaces have property $(G)$ with a modulus which is easily computable in terms of the modulus of uniform convexity.
Proposition 2.8.
CAT(0) spaces have property $(G)$.
Proof.
Let $X$ be a CAT(0) space and $r$, $\varepsilon>0$, $a$, $x$, $y\in X$ with $d(x,a)\leq r$, $d(y,a)\leq r$, $d(x,y)\geq\varepsilon$. By (1), we have that
$d^{2}\left(\frac{x+y}{2},a\right)\leq\frac{1}{2}d^{2}(x,a)+\frac{1}{2}d^{2}(y,a)-\frac{1}{4}d^{2}(x,y)\leq\frac{1}{2}d^{2}(x,a)+\frac{1}{2}d^{2}(y,a)-\frac{\varepsilon^{2}}{4}.$
We do not know (and we leave it as an open problem) whether, generally, $UCW$-hyperbolic spaces have property $(G)$; fortunately, Pinto identified in [18, Proposition 5.4] (see also [22, Proposition 2.5]) the following weaker property of them (which, again, we now codify for the first time) as being enough for the proof to go through.
Definition 2.9.
Let $\psi:(0,\infty)\times(0,\infty)\to(0,\infty)$. We say that a $W$-hyperbolic space $(X,d,W)$ has property $(M)$ with modulus $\psi$ if, for any $r$, $\varepsilon>0$ and any $a$, $x$, $y\in X$ with $d(x,a)\leq r$, $d(y,a)\leq r$, $d(x,y)\geq\varepsilon$, we have that
$d^{2}\left(\frac{x+y}{2},a\right)\leq\max(d^{2}(x,a),d^{2}(y,a))-\psi(r,\varepsilon).$
Thus, the main results of [18] may be expressed by saying that Reich’s theorem and its consequences hold for uniformly smooth hyperbolic spaces having property $(M)$, and in particular property $(G)$.
3 Main results
Let $n\in\mathbb{N}^{*}$ and fix $n$ metric spaces $(X_{1},d_{1}),\ldots,(X_{n},d_{n})$. Put $X:=\prod_{i=1}^{n}X_{i}$ and define $d:X\times X\to X$, by putting, for any $x=(x_{1},\ldots,x_{n})$, $y=(y_{1},\ldots,y_{n})\in X$,
$d(x,y):=\left(\sum_{i=1}^{n}d_{i}^{2}(x_{i},y_{i})\right)^{\frac{1}{2}}.$
It is a classical result that $(X,d)$ is also a metric space.
Now fix $W_{1},\ldots,W_{n}$ such that, for each $i$, $(X_{i},d_{i},W_{i})$ is a $W$-hyperbolic space. Define $W:X^{2}\times[0,1]\to X$ by putting, for any $x=(x_{1},\ldots,x_{n})$, $y=(y_{1},\ldots,y_{n})\in X$ and $\lambda\in[0,1]$, and for any $i$,
$W(x,y,\lambda)_{i}:=W_{i}(x_{i},y_{i},\lambda).$
Proposition 3.1.
$(X,d,W)$ is a $W$-hyperbolic space.
Proof.
We prove $(W1)$. Let $x=(x_{1},\ldots,x_{n})$, $y=(y_{1},\ldots,y_{n})$, $z=(z_{1},\ldots,z_{n})\in X$ and $\lambda\in[0,1]$. We have that
$d(W(x,y,\lambda),z)$ $=\left(\sum_{i=1}^{n}d_{i}^{2}(W(x,y,\lambda)_{i},z_{i})\right)^{\frac{1}{2}}$
$=\left(\sum_{i=1}^{n}d_{i}^{2}(W_{i}(x_{i},y_{i},\lambda),z_{i})\right)^{\frac{1}{2}}$
$\leq\left(\sum_{i=1}^{n}((1-\lambda)d_{i}(x_{i},z_{i})+\lambda d_{i}(y_{i},z_{i}))^{2}\right)^{\frac{1}{2}}$
$=\|((1-\lambda)d_{1}(x_{1},z_{1})+\lambda d_{1}(y_{1},z_{1}),\ldots,(1-\lambda)d_{n}(x_{n},z_{n})+\lambda d_{n}(y_{n},z_{n}))\|$
$=\|(1-\lambda)(d_{1}(x_{1},z_{1}),\ldots,d_{n}(x_{n},z_{n}))+\lambda(d_{1}(y_{1},z_{1}),\ldots,d_{n}(y_{n},z_{n}))\|$
$\leq(1-\lambda)\|(d_{1}(x_{1},z_{1}),\ldots,d_{n}(x_{n},z_{n}))\|+\lambda\|(d_{1}(y_{1},z_{1}),\ldots,d_{n}(y_{n},z_{n}))\|$
$=(1-\lambda)\left(\sum_{i=1}^{n}d_{i}^{2}(x_{i},z_{i})\right)^{\frac{1}{2}}+\lambda\left(\sum_{i=1}^{n}d_{i}^{2}(y_{i},z_{i})\right)^{\frac{1}{2}}$
$=(1-\lambda)d(x,z)+\lambda d(y,z).$
We prove $(W2)$. Let $x=(x_{1},\ldots,x_{n})$, $y=(y_{1},\ldots,y_{n})\in X$ and $\lambda$, $\lambda^{\prime}\in[0,1]$. We have that
$d(W(x,y,\lambda),W(x,y,\lambda^{\prime}))$ $=\left(\sum_{i=1}^{n}d_{i}^{2}(W(x,y,\lambda)_{i},W(x,y,\lambda^{\prime})_{i})\right)^{\frac{1}{2}}$
$=\left(\sum_{i=1}^{n}d_{i}^{2}(W_{i}(x_{i},y_{i},\lambda),W_{i}(x_{i},y_{i},\lambda^{\prime}))\right)^{\frac{1}{2}}$
$=\left(\sum_{i=1}^{n}|\lambda-\lambda^{\prime}|^{2}d_{i}^{2}(x_{i},y_{i})\right)^{\frac{1}{2}}$
$=|\lambda-\lambda^{\prime}|\left(\sum_{i=1}^{n}d_{i}^{2}(x_{i},y_{i})\right)^{\frac{1}{2}}$
$=|\lambda-\lambda^{\prime}|d(x,y).$
$(W3)$ is immediate. We now prove $(W4)$. Let $x=(x_{1},\ldots,x_{n})$, $y=(y_{1},\ldots,y_{n})$, $z=(z_{1},\ldots,z_{n})$, $w=(w_{1},\ldots,w_{n})\in X$ and $\lambda\in[0,1]$. We have that
$d(W(x,y,\lambda),W(z,w,\lambda))$ $=\left(\sum_{i=1}^{n}d_{i}^{2}(W(x,y,\lambda)_{i},W(z,q,\lambda)_{i})\right)^{\frac{1}{2}}$
$=\left(\sum_{i=1}^{n}d_{i}^{2}(W_{i}(x_{i},y_{i},\lambda),W_{i}(z_{i},w_{i},\lambda))\right)^{\frac{1}{2}}$
$\leq\left(\sum_{i=1}^{n}((1-\lambda)d_{i}(x_{i},z_{i})+\lambda d_{i}(y_{i},w_{i}))^{2}\right)^{\frac{1}{2}}$
$=\|((1-\lambda)d_{1}(x_{1},z_{1})+\lambda d_{1}(y_{1},w_{1}),\ldots,(1-\lambda)d_{n}(x_{n},z_{n})+\lambda d_{n}(y_{n},w_{n}))\|$
$=\|(1-\lambda)(d_{1}(x_{1},z_{1}),\ldots,d_{n}(x_{n},z_{n}))+\lambda(d_{1}(y_{1},w_{1}),\ldots,d_{n}(y_{n},w_{n}))\|$
$\leq(1-\lambda)\|(d_{1}(x_{1},z_{1}),\ldots,d_{n}(x_{n},z_{n}))\|+\lambda\|(d_{1}(y_{1},w_{1}),\ldots,d_{n}(y_{n},w_{n}))\|$
$=(1-\lambda)\left(\sum_{i=1}^{n}d_{i}^{2}(x_{i},z_{i})\right)^{\frac{1}{2}}+\lambda\left(\sum_{i=1}^{n}d_{i}^{2}(y_{i},w_{i})\right)^{\frac{1}{2}}$
$=(1-\lambda)d(x,z)+\lambda d(y,w).$
Proposition 3.2.
Assume that, for each $i$, $(X_{i},d_{i},W_{i})$ has property $(G)$. Then $(X,d,W)$ has property $(G)$.
Proof.
For each $i$, let $\psi_{i}$ be a modulus for the property $(G)$ of $(X_{i},d_{i},W_{i})$. Define $\psi:(0,\infty)\times(0,\infty)\to(0,\infty)$ by putting, for any $r$, $\varepsilon>0$,
$\psi(r,\varepsilon):=\min_{i}\ \psi_{i}\left(r,\frac{\varepsilon}{\sqrt{n}}\right).$
We will show that $(X,d,W)$ has property $(G)$ with modulus $\psi$.
Let $r$, $\varepsilon>0$ and $a=(a_{1},\ldots,a_{n})$, $x=(x_{1},\ldots,x_{n})$, $y=(y_{1},\ldots,y_{n})\in X$ with $d(x,a)\leq r$, $d(y,a)\leq r$, $d(x,y)\geq\varepsilon$. It is immediate that, for each $i\in\{1,\ldots,n\}$, $d_{i}(x_{i},a_{i})\leq r$ and $d_{i}(y_{i},a_{i})\leq r$. Assume towards a contradiction that, for every $i\in\{1,\ldots,n\}$, $d_{i}(x_{i},y_{i})<\varepsilon/\sqrt{n}$. Then
$\varepsilon^{2}\leq d^{2}(x,y)=\sum_{i=1}^{n}d_{i}^{2}(x_{i},y_{i})<n\cdot\frac{\varepsilon^{2}}{n}=\varepsilon^{2},$
a contradiction. Thus, there is a $j\in\{1,\ldots,n\}$ such that $d_{j}(x_{j},y_{j})\geq\varepsilon/\sqrt{n}$.
We have, using Proposition 2.2 for the first inequality, that
$\frac{1}{2}d^{2}(x,a)+\frac{1}{2}d^{2}(y,a)-d^{2}\left(\frac{x+y}{2},a\right)$ $=\sum_{i=1}^{n}\left(\frac{d_{i}^{2}(x_{i},a_{i})+d_{i}^{2}(y_{i},a_{i})}{2}-d_{i}^{2}\left(\frac{x_{i}+y_{i}}{2},a_{i}\right)\right)$
$\geq\frac{d_{j}^{2}(x_{j},a_{j})+d_{j}^{2}(y_{j},a_{j})}{2}-d_{j}^{2}\left(\frac{x_{j}+y_{j}}{2},a_{j}\right)$
$\geq\psi_{j}\left(r,\frac{\varepsilon}{\sqrt{n}}\right)$
$\geq\psi(r,\varepsilon),$
which is what we wanted to show. ∎
We leave open the questions of whether property $(M)$ and the property of being $UCW$-hyperbolic are preserved under products.
Now fix $\pi_{1},\ldots,\pi_{n}$ such that, for each $i$, $(X_{i},d_{i},W_{i},\pi_{i})$ is a smooth hyperbolic space. Define $\pi:X^{2}\times X^{2}\to\mathbb{R}$ by putting, for any $x=(x_{1},\ldots,x_{n})$, $y=(y_{1},\ldots,y_{n})$, $u=(u_{1},\ldots,u_{n})$, $v=(v_{1},\ldots,v_{n})\in X$,
$\pi(\overrightarrow{xy},\overrightarrow{uv}):=\sum_{i=1}^{n}\pi_{i}(\overrightarrow{x_{i}y_{i}},\overrightarrow{u_{i}v_{i}}).$
Proposition 3.3.
$(X,d,W,\pi)$ is a smooth hyperbolic space.
Proof.
$(P1)$, $(P2)$ and $(P3)$ are immediate. We now prove $(P4)$. Let $x=(x_{1},\ldots,x_{n})$, $y=(y_{1},\ldots,y_{n})$, $u=(u_{1},\ldots,u_{n})$, $v=(v_{1},\ldots,v_{n})\in X$. We have that
$\pi(\overrightarrow{xy},\overrightarrow{uv})$ $=\sum_{i=1}^{n}\pi_{i}(\overrightarrow{x_{i}y_{i}},\overrightarrow{u_{i}v_{i}})$
$\leq\sum_{i=1}^{n}d_{i}(x_{i},y_{i})d_{i}(u_{i},v_{i})$
$=\langle(d_{1}(x_{1},y_{1}),\ldots,d_{n}(x_{n},y_{n})),(d_{1}(u_{1},v_{1}),\ldots,d_{n}(u_{n},v_{n}))\rangle$
$\leq\|(d_{1}(x_{1},y_{1}),\ldots,d_{n}(x_{n},y_{n}))\|\cdot\|(d_{1}(u_{1},v_{1}),\ldots,d_{n}(u_{n},v_{n}))\|$
$=\left(\sum_{i=1}^{n}d_{i}^{2}(x_{i},y_{i})\right)^{\frac{1}{2}}\cdot\left(\sum_{i=1}^{n}d_{i}^{2}(u_{i},v_{i})\right)^{\frac{1}{2}}$
$=d(x,y)d(u,v).$
We prove $(P5)$. Let $x=(x_{1},\ldots,x_{n})$, $y=(y_{1},\ldots,y_{n})$, $z=(z_{1},\ldots,z_{n})\in X$ and $\lambda\in[0,1]$. We have that
$d^{2}(W(x,y,\lambda),z)$ $=\sum_{i=1}^{n}d_{i}^{2}(W(x,y,\lambda)_{i},z_{i})$
$=\sum_{i=1}^{n}d_{i}^{2}(W_{i}(x_{i},y_{i},\lambda),z_{i})$
$\leq\sum_{i=1}^{n}\left((1-\lambda)^{2}d_{i}^{2}(x_{i},z_{i})+2\lambda\pi_{i}(\overrightarrow{y_{i}z_{i}},\overrightarrow{W_{i}(x_{i},y_{i},\lambda)z_{i}})\right)$
$=\sum_{i=1}^{n}\left((1-\lambda)^{2}d_{i}^{2}(x_{i},z_{i})+2\lambda\pi_{i}(\overrightarrow{y_{i}z_{i}},\overrightarrow{W(x,y,\lambda)_{i}z_{i}})\right)$
$=(1-\lambda)^{2}d^{2}(x,z)+2\lambda\pi(\overrightarrow{yz},\overrightarrow{W(x,y,\lambda)z}).$
Proposition 3.4.
Assume that, for each $i$, $(X_{i},d_{i},W_{i},\pi_{i})$ is a uniformly smooth hyperbolic space. Then $(X,d,W,\pi)$ is a uniformly smooth hyperbolic space.
Proof.
For each $i$, let $\omega_{i}$ be a modulus of uniform continuity for $\pi_{i}$. Define $\omega:(0,\infty)\times(0,\infty)\to(0,\infty)$ by putting, for any $r$, $\varepsilon>0$,
$\omega(r,\varepsilon):=\min_{i}\ \omega_{i}\left(r,\frac{\varepsilon}{\sqrt{n}}\right).$
We will show that $\omega$ is a modulus of uniform continuity for $\pi$.
Let $r$, $\varepsilon>0$ and $a=(a_{1},\ldots,a_{n})$, $u=(u_{1},\ldots,u_{n})$, $v=(v_{1},\ldots,v_{n})$, $x=(x_{1},\ldots,x_{n})$, $y=(y_{1},\ldots,y_{n})\in X$ with $d(u,a)\leq r$, $d(v,a)\leq r$ and $d(u,v)\leq\omega(r,\varepsilon)$. We want to show that
$|\pi(\overrightarrow{xy},\overrightarrow{ua})-\pi(\overrightarrow{xy},\overrightarrow{va})|\leq\varepsilon\cdot d(x,y).$
It is immediate that, for each $i$, $d_{i}(u_{i},a_{i})\leq r$, $d_{i}(v_{i},a_{i})\leq r$ and
$d_{i}(u_{i},v_{i})\leq\omega(r,\varepsilon)\leq\omega_{i}\left(r,\frac{\varepsilon}{\sqrt{n}}\right).$
Thus, for each $i$,
$|\pi_{i}(\overrightarrow{x_{i}y_{i}},\overrightarrow{u_{i}a_{i}})-\pi_{i}(\overrightarrow{x_{i}y_{i}},\overrightarrow{v_{i}a_{i}})|\leq\frac{\varepsilon}{\sqrt{n}}\cdot d_{i}(x_{i},y_{i}).$
We have that
$|\pi(\overrightarrow{xy},\overrightarrow{ua})-\pi(\overrightarrow{xy},\overrightarrow{va})|$ $=\left|\sum_{i=1}^{n}\left(\pi_{i}(\overrightarrow{x_{i}y_{i}},\overrightarrow{u_{i}a_{i}})-\pi_{i}(\overrightarrow{x_{i}y_{i}},\overrightarrow{v_{i}a_{i}})\right)\right|$
$\leq\sum_{i=1}^{n}\left|\pi_{i}(\overrightarrow{x_{i}y_{i}},\overrightarrow{u_{i}a_{i}})-\pi_{i}(\overrightarrow{x_{i}y_{i}},\overrightarrow{v_{i}a_{i}})\right|$
$\leq\frac{\varepsilon}{\sqrt{n}}\sum_{i=1}^{n}d_{i}(x_{i},y_{i})$
$=\frac{\varepsilon}{\sqrt{n}}\cdot\langle(1,\ldots,1),(d_{1}(x_{1},y_{1}),\ldots,d_{n}(x_{n},y_{n}))\rangle$
$\leq\frac{\varepsilon}{\sqrt{n}}\cdot\|(1,\ldots,1)\|\cdot\|(d_{1}(x_{1},y_{1}),\ldots,d_{n}(x_{n},y_{n}))\|$
$=\frac{\varepsilon}{\sqrt{n}}\cdot\sqrt{n}\cdot\left(\sum_{i=1}^{n}d_{i}^{2}(x_{i},y_{i})\right)^{\frac{1}{2}}$
$=\varepsilon\cdot d(x,y).$
Denote by $\mathbb{R}^{2}_{3}$ the Euclidean plane equipped with the $3$-norm. It is known that this is a uniformly convex and uniformly smooth Banach space which is not a Hilbert space. From the above results, it follows that ×$\mathbb{H}\times\mathbb{R}^{2}_{3}$ is a uniformly smooth hyperbolic space that has property $(G)$, and, thus, Reich’s theorem and its consequences hold for it. We will now show that this space is neither a CAT(0) space, nor a convex subset of a normed space (with the canonical convexity structure).
Proposition 3.5.
$\mathbb{R}^{2}_{3}$ is not a CAT(0) space.
Proof.
It is known that a Banach space is a CAT(0) space iff it is Hilbert. ∎
Corollary 3.6.
×$\mathbb{H}\times\mathbb{R}^{2}_{3}$ is not a CAT(0) space.
Proof.
We use the fact that $\mathbb{R}^{2}_{3}$ can be seen as a convex subset of ×$\mathbb{H}\times\mathbb{R}^{2}_{3}$. ∎
Proposition 3.7.
$\mathbb{H}$ is not a convex subset of a normed space.
Proof.
It is immediate that any $W$-hyperbolic space $X$ arising from the canonical convexity structure of a convex subset of a normed space satisfies the following: for any $x$, $y$, $z\in X$ and any $\lambda\in[0,1]$,
$d(W(x,z,\lambda),W(y,z,\lambda))=(1-\lambda)d(x,y).$
We will now exhibit $x$, $y$, $z\in\mathbb{H}$ and $\lambda\in[0,1]$ such that
$d(W(x,z,\lambda),W(y,z,\lambda))<(1-\lambda)d(x,y).$
Take $x:=(0,1)$, $y:=(1,1)$, $z:=(0,2)$, $\lambda:=1/2$. By (3), we have that $W(x,z,\lambda)=(0,\sqrt{2})$ and $W(y,z,\lambda)=(2/3,2\sqrt{5}/3)$. Using (2), we get that
$d(x,y)=\operatorname{arcosh}(3/2)=\ln((3+\sqrt{5})/2)$
and
$d(W(x,z,\lambda),W(y,z,\lambda))=\operatorname{arcosh}(7/(2\sqrt{10}))=\ln(\sqrt{10}/2).$
We have to show that
$\ln(\sqrt{10}/2)<\ln((3+\sqrt{5})/2)/2.$
But this follows from the strict inequality
$(\sqrt{10}/2)^{2}<(3+\sqrt{5})/2,$
which is immediate. ∎
Corollary 3.8.
×$\mathbb{H}\times\mathbb{R}^{2}_{3}$ is not a convex subset of a normed space.
Proof.
We use the fact that $\mathbb{H}$ can be seen as a convex subset of ×$\mathbb{H}\times\mathbb{R}^{2}_{3}$. ∎
4 Acknowledgements
This work has benefited from discussions with Pedro Pinto, the principal investigator of the German Science Foundation (DFG) project Ref. PI 2070/1-1, Projektnummer 549333475, in which this problem is listed as an objective and I am listed as a collaborator.